Two cars leave Denver at the same time and travel in opposite direction. One car travels 10m/h faster than the other car . the cars are 500 miles apart in 5 h. how fast is each car driving.
Found 2 solutions by checkley77, TR Tutoring:
Answer by checkley77(12569)

(x+x+10)5=500
(2x+10)5=500
10x+50=500
10x=500-50
10x=450
x=450/45
x=45 mph. for the slower car.
45+10=55 mph. for the faster car.
Proof:
(45+55)5=500
100*5=500
500=500

Answer by TR Tutoring(2)

This problem is a copyright violation of the curriculum used by Connections Academy, an online charter school. Students who use this problem and it's answers on this website are cheating on their math test. Please remove this problem from your website.